a basketball player has a vertical leap of about 36in.If the equation d=48t^2 describes his hang timet is a function of his verticalleap d,what is his hang time?
Answer by stanbon(75887)

a basketball player has a vertical leap of about 36in.If the equation d=48t^2 describes his hang timet is a function of his verticalleap d,what is his hang time?
----------------------------------------
36 = 48t^2
t^2 = 3/4
t = (1/2)sqrt(3) seconds
==============================
Cheers,
Stan H.